On August 1, 2013, Ace Corporation accepted a note receivable in place of an outstanding accounts receivable in the amount of $123,965. The note is due in 90 days and has an interest rate of 8%. What would be the appropriate journal entry to record the receipt of cash at the maturity date?

Select one:

a.

Cash

123,965.00

             Notes Receivable

123,965.00

b.

Notes Receivable

123,965.00

Interest Revenue

2,479.30

              Cash

126,444.30

c.

Cash

126,444.30

          Interest Revenue

2,479.30

          Notes Receivable

123,965.00

d.

Notes Receivable

123,965.00

             Accounts Receivable

123,965.00
Answer #1

Answer

  • Correct Answer = Option ‘c’

Cash

126,444.30

          Interest Revenue

2,479.30

          Notes Receivable

123,965.00

Interest Revenue will be for 3 months (or 90 days) = $ 123965 x 8% x 3/12 = $ 2479.30

Cash received = 123965 + 2479.30 = $ 126444